Can someone please help me with this problem:
Find the exact interest. Use 365 days in a year, and use the exact number of days in a month. Round to the nearest cent, if necessary.
A loan of $76,000 at 11% made on Feb 29 and due on June 30.

Answer by scott8148(6628)

March -- 31 days
April -- 30 days
May -- 31 days
June -- 30 days

total days -- 122

interest = 76000 * 11% * 122/365 = 760 * 11 * 122/365 = 2794.30